SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 2 September 2004

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Thursday, September 2, 2004

image SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.   Today was a "half-duty day" in view of tomorrow's EVA load. 

EVA-11 preparations and systems reconfigurations proceeded nominally, both onboard and on the ground.  U.S. & Russian EVA support teams report ready to conduct the spacewalk.

The crew tagged up with TsUP specialists for a final review of the timeline and procedures.  There were no questions.   [The final timeline was slightly "tweaked" from yesterday's version (by modifying the handrail-to-handrail translation path from the DC1's EVA ladder to the FGB as per crew recommendation, and narrowing down the designated Safety Zone on the Service Module (SM) to the circumferential handrails on the conical section of the SM Working Compartment (RO).]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

FE/SO Fincke prepared cabin systems for the isolation of the Russian segment (RS) from the U.S. segment (USOS) during the EVA, by transferring SODF (Systems Operations Data File) items and the CCPK (Crew Contamination Protection Kit) from the USOS to the SM.   [SODF items include hardcopies of the SODF Warning book, POC book and Medical Checklist, as well as portions of the CD library.]

In addition, the FE printed out USOS unmanned configuration procedures on the Lab printer for transferal to the SM.

Fincke also relocated a PCS (portable computer system) 760XD laptop from the LAB RWS (robotic work station) to the FGB.

Other station preparations by Mike for the 5h 49m spacewalk included:

  • Configuring the EPS (electrical power system) as required for USOS/RS module isolation for EVA (verifying that all U.S. GLAs (general luminaire assemblies) manual switches, except for the failed ones, are in the On position and lights illuminated);
  • Reconnecting the UOP (utility outlet panel) bypass power cable at both Lab and Cupola RWS DCPs (display & control panels);
  • Setting up two Sony PD100 camcorders in the Lab & Node for situational awareness during the spacewalk; and
  • Installing & testing the freshly charged NiMH (nickel metal hydride) batteries in the EHIP (EMU helmet interchangeable portable) lights, borrowed from U.S. EMUs and jerry-rigged on the Orlans.

Tasks worked by Gennady Padalka preparatory to the EVA included:

  • Filling the U.S. DIDBs (disposable in-suit drink bags) with water and installing them in the Orlans;
  • Equipping both Orlans (in pocket on left calf) with a "Pille-MKS" radiation sensor.   [Two sensors, A0309 &A0310, were removed from their exposure locations in the RS after recording their dose measurements; a third, A0307, was placed in the SM for background readings.  Also, tomorrow when donning his gear, Padalka will place his ID-3 personal dosimeter, normally worn on the flight suit, into the chest pocket in his Orlan's lining (near the DIDB) and later return it to the flight suit.]
  • Removing the Kriogem-03M refrigerator equipment from the DC1 and transferring it to the SM Transfer Tunnel (PrK) for temporary storage.   [The cooler currently contains the Japanese JCF-01V crystallizer unit of the GCF (Granada Crystallization Facility) and LUCH-2 biological crystallization hardware.]; and
  • Powering down the new Sputnik-SM "Kenwood D700" amateur radio station to protect the spacewalkers from inadvertent antenna RF emission.

Mike Fincke performed the regular once-a-week maintenance reboot on the operational PCS laptops.

The CDR transferred the accumulated data files of his BIO-5 Rasteniya-2 ("Plants-2") plant growth experiment to the Packet laptop for subsequent downlink to the ground via Regul comm.

Gennady also completed another periodic replenishing of the Elektron's water supply for electrolysis, filling the KOV thermal loops' EDV container with purified (deionized) water from the BKO multifiltration/purification column unit.   [Elektron is currently operating nominally but will be shut down tomorrow 30 min prior to Orlan donning.  Also to be turned off in the RS for the duration of the spacewalk: SKV-1 air conditioner, IP-1 air flow sensors, some heaters, VHF channel 1 (except over RGS when it's On), automatic monitoring of DC1 smoke detectors, the GTS (Global Timing System) 400MHz transmitter shortly before DC1 hatch opening, etc.]

The FE took the periodic (weekly) reading of the cabin air's current CO2 partial pressure in the SM and Lab, using the U.S. CDMK (CO2 monitor kit), for calldown (along with the battery status) for use in trending analyses.

Padalka conducted today's routine inspection of the SM's SOZh life support system (including replacement of ASU toilet facility inserts), and Mike handled the routine status checkup of the autonomous PCG-STES010 (Protein Crystal Growth-Single Locker Thermal Enclosure System) payload in the Lab, ahead one day from the regular schedule because of the EVA.

The crew completed its daily physical exercise on TVIS treadmill, RED resistive exerciser, and VELO stationary bike with load trainer.

The CDR once more broke out the "Urolux" equipment, setting it up for the Russian PZE MO-9 biochemical urine test tomorrow, a standard requirement before and after Orlan-suited activities.   [The MO-9 analysis uses the sophisticated in-vitro diagnostic apparatus "Urolux" developed originally for the Mir program.]

Mike's load calibration of the new SchRED (Schwinn resistive exercise device) flexpack canisters yesterday was reviewed by the ground and found to "look great", matching pre-flight data generally within 1-2 lbs.

Yesterday's station atmosphere repress with air from Progress 15 tankage went smoothly, raising total pressure by ~25 mmHg/Torr to 764.

Yesterday, the ground turned on shell heaters in the PMA-1 (Pressurized Mating Adapter 1) and Node to provide the preheat condition required before the planned preemptive power downs tomorrow.  Today, the crew also manually turned on the TCS (thermal control system) shell heater in the U.S. Airlock for that purpose.

For his next "Saturday Science" program SO Fincke has selected the HEAT experiment.  This requires the MSG (Microgravity Science Glovebox) in the Lab, and European MSG specialists will be on standby.   [The European HEAT was one of three experiments of André Kuipers' VC-6 "Delta" science program that were performed in the USOS.  Installed in the MSG, it was to investigate heat transfer performance of grooved heat, but the experiment failed to work properly during the VC-6 joint crew period.  Mike now volunteered to run it again.]

The FE was successful yesterday in "ghosting" (loading) the two ePCS (early Portable Computer System) hard drives with the new software version R2.   [This enables the crew to restart the C&C MDMs (command & control computers) from ePCS laptop and allows them to monitor some of the Lab systems in the unlikely event of all three C&Cs failing.  The previous ePCS version (e5A) did not have the capability of restarting the C&Cs when the resident special fail-safe software package nicknamed "Mighty Mouse" is unable to recover the three MDMs.]

Several times today, when over RGS (Russian ground sites), MCC-Houston tested transmitting U.S. contingency telemetry through RGS and S-band.  These checkouts have already been conducted successfully last Monday (8/30) but were repeated because of changes made since then for the software uploading on 8/31.

The long-awaited task of evacuating the Lab window's inter-pane space ("Volume D") and installing the new U-jumper flexhose and its protective cover is now scheduled for 9/17 (or so).

Sleep cycle will begin tonight 15 min later than usual, for a ten-hour period of night rest.  Wakeup is at 3:45am tomorrow morning.

Since 9:38pm last night, the ISS is again flying in earth-oriented LVLH (local vertical/local horizontal) attitude (yaw: 350 deg, pitch: 352.8 deg, roll: 0 deg), in preparation for EVA-11.

Update on Consumables:  Onboard resources continue to be monitored carefully.  All supplies (water, N2, O2, food, propellant) are tracking close to predictions.   [Water is just below planning due to lower condensate collection, prop quantity is above because of postponed reboosts and orbital decay having been less than projected.  Nitrogen: below planning due to EVA represses after EVA-9b and -10.  Food: six food containers (3 US, 3 Russian) manifested on Soyuz 9S.]

Update on SM Printer Failure: The malfunction has been found to be similar to an SM printer failure during Expedition 5, which could not be recovered as yet.  Therefore, there will be no further troubleshooting for the crew and instead the ground is working to get another printer sent up for Expedition 10.  Also, there still is the option to move the Lab printer to the SM.

Update on Soyuz 9S:  Launch of CDR Leroy Chiao and FE Salizhan Sharipov continues to be set for 10/9.  Their Soyuz TMA-5 spacecraft is the first with two new features that are welcome improvements of the reliable old crew transport: two additional forward-pointing braking thrusters (#27, #28) besides the two engines (#17, #18) already near the Orbital Module's docking ring; and a thermo-electric cooler for the Descent Module's Hydrogen Peroxide tankage, to extend the life of the H2O2 which tends to deteriorate in time to H2O and O.  (H2O2 is one of the most powerful oxidizers known -- stronger than chlorine, chlorine dioxide, and potassium permanganate, but it has been (and still is, until certification) limiting Soyuz' orbital stay time).   Major upcoming events:

  • EVA-11 -- 9/3 (hatch open 12:50pm EDT);
  • Soyuz 9S launch -- 10/9;
  • Soyuz 9S dock -- 10/11;
  • Soyuz 8S undock/land -- 10/19;
  • Soyuz 9S relocate -- 11/18;
  • Progress 16P launch -- 11/24.

EVA-11 (tomorrow, 9/3) -- Main timeline activities (all times Eastern Daylight):

  • 11:09am -- isolate DC1 airlock from ISS/RS interior;
  • 12:50pm -- open DC1 EV1 hatch;
  • 1:01pm -- egress DC1;
  • 1:27pm -- translate to FGB & secure sealed container (GK) with spare flow regulator valve (RRZh);
  • 1:34pm -- remove old RRZh from FGB external coolant loop #1 & install new one;
  • 1:53pm -- install fairleads (spiral shaped tether holders) on FGB handrails;
  • 1:56-2:26pm -- night pass (eclipse); conduct "immobility" experiment
  • 2:26pm -- install fairleads;
  • 2:38pm -- return old RRZh valve in airtight GK container to DC1;
  • 3:06pm -- take photos of BKDO contamination monitoring unit on DC1 and MPAC/SEED units on SM;
  • 3:26pm -- translate to rear of SM;
  • 3:28-3:58pm -- night pass (eclipse); crew rest;
  • 3:59pm -- install three WAL antennas for the ATV;
  • 4:59-5:30pm -- night pass (eclipse); crew rest;
  • 5:30pm -- install three WAL antennas for the ATV & remove covers from all 5 ATV antennas;
  • 6:12pm -- translate back to DC1;
  • 6:31pm -- install handrail limiters;
  • 6:37pm -- ingress DC1;
  • 6:45pm -- close EV hatch 1.

U.S. CMGs will have attitude control throughout the EVA.  SM thrusters will be inhibited at 2:30pm, prior to the crew's translation to the SM and uninhibited at 6:30pm on the crew's return to the DC1.  CMG momentum saturation is not expected to occur, but if desaturation should indeed become necessary the crew will get out of the "keep out" zones on the SM, with the station in free drift (for up to one orbit) before thrusters are "armed".  In case of comm loss (LOS) during orbital day, the crew will proceed with task execution until the next eclipse (night pass), then retreat to the safe zone (SM conical section).

More information and background material on EVA 4 can be found at the Expedition 9 EVA 4 Reference Page

Expedition 9 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today's CEO photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Internal waves, Azores Is - Mid Atlantic (continued clearing in the north Atlantic provides good conditions for photography of internal waves.  The sunglint point was to the left of track.  Some interesting features may have been visible in the westernmost Azores), Hurricane Frances (Dynamic Event.  Frances was hammering the Bahamas at the time of this pass.  Looking to the right of track for the eye), River dynamics, Beni River, Bolivia (clear weather and a nadir pass provide an excellent mapping opportunity for the Beni River.  The fan expands to the west. The ground is particularly interested in mapping of the current and past river courses), and Hurricane Howard (Dynamic Event.  Howard was predicted to become a Category 2 storm by the time of this near-nadir overpass.  The predicted course had Howard moving out to sea, but photography of the eye and banding (to the right of track) will be useful).

CEO images can be viewed at these websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 1:30pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Sampling Mode, following Zero Calibration.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On. SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) -- 767; temperature (deg C) -- 26.5; ppO2 (mmHg) -- 168.7; ppCO2 (mmHg) -- 3.8.
  • SM Transfer Compartment:  Pressure (mmHg) -- 761; temperature (deg C) -- 21.6.
  • FGB Cabin:  Pressure (mmHg) -- 756; temperature (deg C) -- 23.3.
  • Node:  Pressure (mmHg) -- 765.5; temperature (deg C) -- 27.4 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab:  Pressure (mmHg) -- 767.4; temperature (deg C) -- 23.4; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) -- 767.5; temperature (deg C) -- 24.4; shell heater temp (deg C) -- n/a, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • (n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, "sun slicer", i.e., drag reduction-biased by 47 deg angle (2B: +47, 4B: -47).
  • SM batteries:  All batteries (8) are on line in "Partial Charge" mode.  
  • FGB batteries:  Battery #3 is off line.  All other batteries (5) are on line in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System (PS):

  • Total propellant load available: 4239 kg (9345 lb) as of 8/26/04;  [SM(552) + FGB(3407) + Progress M(280)].  (Capabilities: SM -- 860 kg; FGB -- 6120 kg).

Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2's RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- U.S. SIGI-1 (GPS)
  • Angular rate source -- RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: 180.5 deg, pitch: -6.9 deg., roll: 0 deg]), with CMG TA (thruster assist) until 9/2 for EVA-11, then back to XPOP until next reboost (9/22).

Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit  (as of this morning, 6:43am EDT [= epoch]):

  • Mean altitude -- 361.0 km
  • Apogee height -- 364.8 km
  • Perigee height -- 357.3 km
  • Period -- 91.76 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.0005542
  • Solar Beta Angle -- 42.3 deg (magnitude increasing)
  • Orbits per 24-hr. day -- 15.69
  • Mean altitude loss in last 24 hours -- 100 m
  • Revolutions since FGB/Zarya launch (Nov. '98)  -- 33045

ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.